The metropolitan area of Los Angeles–Long Beach, California, showed the largest increase in the number of jobs between 1988 and 2003. The number of jobs increased from 5,152,500 to 6,286,100. What was the geometric mean expected yearly rate of increase? (Round your answer to 2 decimal places.)

  Geometric mean  %
Answer #1

Given-

Vi = initial value = 5,152,500

Vf = final value = 6,286,100

N = number of years = 2003 - 1988 = 15

The formula for calculating the geometric mean (also known as the exponential growth rate) is:

R = (Vf / Vi)1 / N

= (6,286,100 / 5,152,500)1 / 15

= 1.22 1 / 15

= 1.013345

=> The geometric mean expected yearly rate of increase is 1.013345 - 1 = 0.013345

=> 1.33 % growth rate per year